# Artificial Intelligence in Image-based Cardiovascular Disease Analysis: A Comprehensive Survey and Future Outlook

### v3

@@ 2.4 Cardiac Ultrasound (Echocardiography)
Cardiac ultrasound imaging is a fundamental tool in cardiovascular assessment, employing sound waves to produce real-time images of the heart [55, 56]. These images reveal critical details about the heart’s size, shape, function, and blood flow patterns [57]. Echocardiography’s non-invasive approach, combined with the absence of ionizing radiation and its capability for immediate imaging, makes it a preferred choice for initial diagnosis in a range of cardiac conditions [58]. It is crucial in identifying and managing various heart diseases, such as valvular heart disease, cardiomyopathies, and congenital heart disease [59, 21].
For the vessels, Intravascular Ultrasound (IVUS) is a medical imaging methodology used in cardiology to visualize the inside of the heart’s coronary arteries from within the artery itself [60]. It involves using a specially designed catheter with a miniaturized ultrasound probe attached to its distal end. The catheter is threaded through the coronary vasculature to the area of interest, and ultrasound is used to produce detailed images of the coronary arteries’ interior walls. [61].
@@ 2.5 Nuclear Cardiology
Nuclear imaging in cardiology utilizes small amounts of radioactive tracers injected into the patient’s bloodstream. These tracers emit gamma rays, captured by specialized cameras to generate images of the heart’s structure and function [62]. Nuclear cardiology, including single-photon emission computed tomography (SPECT) and positron emission tomography (PET), have been a key non-invasive imaging modality for patients with known or suspected cardiovascular disease.
Cardiac Single-Photon Emission Computed Tomography (SPECT): SPECT is a vital diagnostic tool for evaluating heart function and blood flow. As a non-invasive imaging technique, it offers essential information regarding the heart’s structure, blood supply, and overall health. It plays a significant role in diagnosing, managing, and monitoring various cardiovascular conditions, contributing to better patient care and treatment planning [63].
Positron Emission Tomography (PET): PET is a highly advanced and non-invasive imaging technique used in the evaluation of heart diseases. It utilizes a small amount of radioactive material, a PET scanner, and a computer to evaluate the function and metabolism of the heart. This technique stands out for its exceptional accuracy in detecting coronary artery disease, assessing myocardial perfusion (blood flow to the heart muscle), and evaluating heart function. Cardiac PET is particularly effective in identifying areas of reduced blood flow, differentiating between viable and non-viable heart muscle, and diagnosing various cardiac conditions. Its high sensitivity and specificity make it a valuable tool in the management of heart diseases [64].
Figure 3: Examples of recent representative AI models for CVD Analysis. Top (Non-vessel). Examples of Networks and Tasks: Convolutional neural network (CNN) [65], Transformer [66], Fully convolutional networks (FCN) [67], U-Net [68], Autoencoders (AE) and variational autoencoders (VAE) [69], Generative adversarial networks (GAN) [70], and Contrastive Language Image Pre-training (CLIP) [71]. Examples of Applications: Calcium Scoring [72], Image Registration [73], Whole Heart Segmentation [74], Image Synthesis [75], Tagged MRI to Cine MRI Transform [76], and Report Generation [14]. Bottom (Vessel). Examples of Networks and Tasks: Diffusion Model [77], CNN+++Recurrent Neural Networks (RNN) [78], Tree-structured RNN [79], Graph Neural Network (GNN) [80], and Reinforcement Learninging (RL) [81]. Examples of Applications: Self-supervised Vessel Segmentation [82], Vessel Stenosis Detection [83], Vessel Segmentation [79], Vessel Labeling [84] and Guidewire Navigation [85].
@@ 2.6 Cardiac Magnetic Resonance Imaging (MRI)
Cardiac MRI is a non-invasive imaging technique that uses a powerful magnetic field and radio waves to produce detailed images of the heart, offering crucial insights into its structure, function, and blood flow [86, 87]. We discuss several MRI modalities commonly used in cardiac imaging.
@@
Cine MRI: Cine MRI captures a series of images across the cardiac cycle to visualize the beating heart in real time, allowing evaluation of heart structure, ventricular size, wall motion, and ejection fraction. It’s essential in detecting conditions like ventricular hypertrophy, myocardial infarction, and various valvular disorders [88, 89].
@@
Cardiac tagging magnetic resonance imaging (t-MRI). Cardiac tagging magnetic resonance imaging (t-MRI), also known as myocardial tagging or t-MRI, offers a unique approach to evaluating myocardial movement and deformation. During imaging, a grid-like pattern or ”tags” are superimposed on the heart muscle, revealing intricate details about myocardial function. These tags change shape as the heart beats, providing valuable insights into both global and regional cardiac function. Regarded as the gold standard for assessing regional myocardial deformation and strain, cardiac tagging is pivotal in the diagnosis, management, and research of various heart conditions, particularly in diseases such as Ischemic Heart Disease and Dilated Cardiomyopathy [25].
@@
Late Gadolinium Enhancement (LGE) MRI. LGE MRI is a crucial imaging technique that enhances the visibility of myocardial damage or fibrosis by using a gadolinium-based contrast agent. It plays a key role in evaluating myocardial viability and identifying pathological tissues, such as scarred areas. This method is particularly effective in diagnosing and assessing various cardiac conditions, including myocardial infarction, myocarditis, and different types of cardiomyopathies. LGE MRI is also increasingly recognized as a valuable tool for assessing scar tissues in patients with atrial fibrillation (AF), offering an advanced alternative for detailed cardiac evaluation [20].
@@
T2-weighted MRI Images: T2-weighted MRI stands out for its ability to provide enhanced contrast by distinguishing tissues based on water content and physiological state. This feature is particularly valuable in identifying myocardial abnormalities, such as inflammation, edema, or ischemia. T2-weighted MRI is especially effective in detecting myocardial edema caused by inflammation or acute ischemia, offering critical insights for diagnosis and treatment [90].
@@
Myocardial Perfusion MRI: This imaging modality is instrumental in assessing blood flow to the heart muscle, crucial for identifying regions with insufficient blood supply. It involves injecting a contrast agent and capturing images as the contrast first passes through the heart. Myocardial perfusion MRI is essential for diagnosing and managing conditions like coronary artery disease and ischemic heart disease [91, 92].
@@
Diffusion-Weighted Imaging (DWI): DWI provides a unique perspective by evaluating the diffusion of water molecules within tissues, revealing information about tissue microstructure. This technique is particularly useful for assessing areas of acute or chronic myocardial ischemia, as alterations in water diffusion properties can indicate tissue damage [93].
@@
4D Flow MRI: 4D Flow MRI is a cutting-edge technique that visualizes the velocity and direction of blood flow in three dimensions over time. This comprehensive approach to blood flow analysis is crucial for assessing abnormalities in blood vessels, such as aortic aneurysms, aortic dissections, and congenital heart defects [94].
In summary, various imaging modalities are specialized to detect modality-specific features critical to their clinical use for CVD diagnostics [95, 96, 97, 98]. For example, CT scans are utilized to detect characteristics such as density and tissue contrast, which are important for identifying calcifications, while MRI scans concentrate on identifying features related to soft tissue contrast and motion.
Figure 4: Overview of the Cardiac Image Analysis Pipeline and Examples. Top: Non-vessel. (a): Cardiac segmentation [99], End-diastole (ED) and End-systole (ES) phases detection and Surface Model [100]. (b): Cardiac Motion Tracking [100]. Myocardial blood flow (MBF) maps [101]. Middle: Medical images serve as inputs for AI models in the CVD diagnosis pipeline. Bottom: Vessel. (c): Segmentation and quantification of vessel structures, with a focus on the Aorta [102]. Coronary Artery Segmentation [103], centerline extraction using segmentation masks [103], Coronary Artery Branch Labeling based on centerline structures [84], and detection of Coronary Artery Stenosis [104]. (d): Evaluation of vessel function. Arota flow Speed with 3D streamlined visualization[105]. Interpretation of CT-FFR results showing significant hemodynamic impact [34], and Coronary Microvascular [106]. Vessel abbreviations: LAD = left anterior descending; LCX = left circumflex; RCA = right coronary artery. Physiological measurements: CFR = coronary flow reserve; FFR = fractional flow reserve; IMR = index of microvascular resistance.
@@ 3 AI in Individual Imaging CVD Analysis
In this section, we focus on AI pipelines for individual CVD analysis, targeting both non-vessel and vessel cardiac structures. Our aim is to outline prevalent cardiac diseases and their analysis using AI technologies in cardiac imaging. We provide an overview of key diseases, the data modalities used for their study, and recent AI methodologies in this area. The application of some representative AI models in cardiac imaging is depicted in Figure 3, offering insights into the integration of AI in cardiac health and disease management.
3.1 AI Methods and Pipelines
@@
Figure 4 illustrates that cardiac image segmentation plays a crucial role in numerous clinical applications [107, 108]. This process is key to identifying and analyzing essential anatomical structures [109], as it divides the image into regions with distinct semantic meanings and facilitates the extraction of various quantitative metrics. In the context of non-vessel structures, segmentation aids in calculating important metrics such as the volumes and ejection fractions of the left and right ventricles (LV and RV), atrial sizes, myocardial mass, and wall thickness [17]. For vessel structures, segmentation is pivotal for analyzing aspects of coronary arteries, including centerline extraction [110], plaque identification, and other vascular features. Such detailed segmentation provides invaluable insights for both the diagnosis and treatment planning processes.
The U-Net architecture, introduced by [68], is a prominent model in medical image segmentation. Its popularity stems from its flexibility, efficient design, and proven success across various imaging modalities [111]. U-Net comprises two main components: the Encoder and the Decoder. The Encoder path captures semantic and contextual features through downsampling and convolutional blocks, while the Decoder path uses convolutional and upsampling blocks to gradually increase spatial resolution and facilitate pixel-wise classification. U-Net’s defining feature is its skip connections, which bridge each level of the Encoder to the corresponding Decoder level, ensuring the transfer of high-resolution contextual details. This feature promotes the combination of low-level details with high-context information for precise localization. Furthermore, U-Net’s architecture easily integrates with Recurrent Neural Network (RNN) structures. An example of this integration is the use of ConvLSTM [112] by Azad et al. [113], which nonlinearly combines features from both the encoder and the preceding decoder layers. Since its inception, U-Net has established itself as a benchmark in medical image segmentation, inspiring various advanced derivatives that enhance its original design. For a comprehensive review of U-Net variants, see [114]. In addition to U-Net, Generative Adversarial Networks (GANs) also play a significant role in medical image segmentation [115, 116, 117, 118]. More recently, denoising diffusion models, a notable subset of generative models, have recently attracted significant attention in the field of deep learning. They have demonstrated remarkable utility across a wide range of applications, especially in the enhancement of medical image segmentation [119]. These models have demonstrated their effectiveness in generating high-quality segmented images, further expanding the possibilities in medical imaging analysis [120].
@@
As illustrated in the right panel of Figure 4, following segmentation, the AI system engages in feature extraction for precise quantification. This stage involves measuring physical heart parameters like chamber volumes, wall thickness, and myocardial mass [109]. Utilizing the segmented structures and the features derived, AI models then conduct functional analysis. This encompasses the calculation of ejection fraction, wall motion, and blood flow dynamics, crucial indicators of the heart’s efficiency and overall functioning [109]. By integrating segmented images, extracted features, and quantification metrics, the AI system forms a comprehensive assessment of cardiac structure and function, invaluable for diagnosing cardiac conditions and formulating treatment plans. The details of methods in Figure 4 are introduced in Section 3.2 for non-vessel structures and Section 3.3 vessel structures. For clinical utility, the AI pipeline presents results in an easily interpretable format, often through 3D models or quantitative maps, compiled into reports for cardiologists’ evaluation. Additionally, it incorporates clinician feedback to refine models, enhancing accuracy and adaptability to new data and insights. This AI-driven approach in cardiology transforms complex imaging data into practical medical information, potentially revolutionizing the diagnosis and treatment of heart diseases.
@@
Image registration plays a critical role in analyzing heart structures and diseases, especially when merging data from different imaging modalities or tracking disease progression over time [121]. Accurate registration is essential for mapping structural changes precisely and is indispensable in longitudinal studies, treatment planning, and assessing therapeutic efficacy [122]. The task of cardiac image registration is particularly challenging due to the significant variability in cardiac shape and motion among individuals [123]. It often involves aligning heart images captured at different times or under varying conditions, such as different phases of the cardiac cycle or using diverse imaging techniques like CT, MRI, or ultrasound [124]. Given the dynamic nature of the heart and the influence of respiratory movements, non-rigid registration is crucial for cardiac images [73]. This approach is key to understanding disease progression or treatment effects by comparing temporal image sequences. Advancements in AI models are vital for enhancing registration accuracy and efficiency. The development of real-time registration systems is particularly significant for their potential application in surgical or interventional settings [81], promising to significantly improve patient outcomes and procedural precision in cardiac care [125].

### v4

@@ 2.4 Cardiac Ultrasound (Echocardiography)
Cardiac ultrasound imaging is a fundamental tool in cardiovascular assessment, employing sound waves to produce real-time images of the heart [zhou2021artificial, liu2023deep]. These images reveal critical details about the heart’s size, shape, function, and blood flow patterns [aly2021cardiac]. Echocardiography’s non-invasive approach, combined with the absence of ionizing radiation and its capability for immediate imaging, makes it a preferred choice for initial diagnosis in a range of cardiac conditions [lu2023ultrafast]. It is crucial in identifying and managing various heart diseases, such as valvular heart disease, cardiomyopathies, and congenital heart disease [ghorbani2020deep, jone2022artificial]. For the vessels, Intravascular Ultrasound (IVUS) is a medical imaging methodology used in cardiology to visualize the inside of the heart’s coronary arteries from within the artery itself [xu2020fundamentals]. The catheter is threaded through the coronary vasculature to the area of interest, and ultrasound is used to produce detailed images of the coronary arteries’ interior walls [de2002intravascular].
@@ 2.5 Nuclear Cardiology
Nuclear imaging in cardiology utilizes small amounts of radioactive tracers injected into the patient’s bloodstream. These tracers emit gamma rays, captured by specialized cameras to generate images of the heart’s structure and function [tamaki2024current]. Nuclear cardiology, including single-photon emission computed tomography (SPECT) and positron emission tomography (PET), have been a key non-invasive imaging modality for patients with known or suspected cardiovascular disease.
2.5.1 Cardiac Single-Photon Emission Computed Tomography (SPECT)
SPECT is a nuclear imaging technique that provides three-dimensional images of blood flow to organs and tissues, widely used in the assessment of heart diseases. In cardiac applications, SPECT helps evaluate myocardial perfusion—how well blood flows through the heart muscle—by allowing clinicians to see areas with reduced blood flow, which may indicate CAD or previous heart damage [apostolopoulos2023deep].
2.5.2 Positron Emission Tomography (PET)
PET is a highly advanced and non-invasive imaging technique used in the evaluation of heart diseases. It utilizes a small amount of radioactive material, a PET scanner, and a computer to evaluate the function and metabolism of the heart. This technique stands out for its exceptional accuracy in detecting coronary artery disease, assessing myocardial perfusion (blood flow to the heart muscle), and evaluating heart function. Cardiac PET is particularly effective in identifying areas of reduced blood flow, differentiating between viable and non-viable heart muscle, and diagnosing various cardiac conditions. Its high sensitivity and specificity make it a valuable tool in the management of heart diseases [slomka2021quantitative].
@@ 2.6 Cardiac Magnetic Resonance Imaging (MRI)
Cardiac MRI is a non-invasive imaging technique that uses a powerful magnetic field and radio waves to produce detailed images of the heart, offering crucial insights into its structure, function, and blood flow [saeed2015cardiac, bello2019deep, wang2024screening]. We discuss several MRI modalities commonly used in cardiac imaging.
2.6.1 Cine MRI
@@
Cine MRI captures a series of images across the cardiac cycle to visualize the beating heart in real time, allowing evaluation of heart structure, ventricular size, wall motion, and ejection fraction. It’s essential in detecting conditions like ventricular hypertrophy, myocardial infarction, and various valvular disorders [kustner2020cinenet, campello2021multi].
2.6.2 Cardiac tagging magnetic resonance imaging (t-MRI)
@@
Cardiac tagging magnetic resonance imaging (t-MRI), also known as myocardial tagging or t-MRI, offers a unique approach to evaluating myocardial movement and deformation. During imaging, a grid-like pattern or ”tags” are superimposed on the heart muscle, revealing intricate details about myocardial function. These tags change shape as the heart beats, providing valuable insights into both global and regional cardiac function. Regarded as the gold standard for assessing regional myocardial deformation and strain, cardiac tagging is pivotal in the diagnosis, management, and research of various heart conditions, particularly in diseases such as Ischemic Heart Disease and Dilated Cardiomyopathy [ye2021deeptag].
2.6.3 Late Gadolinium Enhancement (LGE) MRI
@@
LGE MRI is a crucial imaging technique that enhances the visibility of myocardial damage or fibrosis by using a gadolinium-based contrast agent. It plays a key role in evaluating myocardial viability and identifying pathological tissues, such as scarred areas. This method is particularly effective in diagnosing and assessing various cardiac conditions, including myocardial infarction, myocarditis, and different types of cardiomyopathies. LGE MRI is also increasingly recognized as a valuable tool for assessing scar tissues in patients with atrial fibrillation (AF), offering an advanced alternative for detailed cardiac evaluation [li2022medical].
2.6.4 T2-weighted MRI Images
@@
T2-weighted MRI stands out for its ability to provide enhanced contrast by distinguishing tissues based on water content and physiological state. This feature is particularly valuable in identifying myocardial abnormalities, such as inflammation, edema, or ischemia. T2-weighted MRI is especially effective in detecting myocardial edema caused by inflammation or acute ischemia, offering critical insights for diagnosis and treatment [ren2022comparison].
2.6.5 Myocardial Perfusion MRI
@@
This imaging modality is instrumental in assessing blood flow to the heart muscle, crucial for identifying regions with insufficient blood supply. It involves injecting a contrast agent and capturing images as the contrast first passes through the heart. Myocardial perfusion MRI is essential for diagnosing and managing conditions like coronary artery disease and ischemic heart disease [xue2020automated, alskaf2022deep].
2.6.6 Diffusion-Weighted Imaging (DWI)
@@
DWI provides a unique perspective by evaluating the diffusion of water molecules within tissues, revealing information about tissue microstructure. This technique is particularly useful for assessing areas of acute or chronic myocardial ischemia, as alterations in water diffusion properties can indicate tissue damage [ueda2022deep].
2.6.7 4D Flow MRI
@@
4D Flow MRI is a cutting-edge technique that visualizes the velocity and direction of blood flow in three dimensions over time. This comprehensive approach to blood flow analysis is crucial for assessing abnormalities in blood vessels, such as aortic aneurysms, aortic dissections, and congenital heart defects [berhane2022deep].
@@ 3 AI in Individual Imaging CVD Analysis
In this section, we focus on AI pipelines for individual CVD analysis across both non-vessel and vessel cardiac structures. Our goal is to summarize key cardiac diseases, the imaging modalities used to evaluate them, and the AI methodologies that support their diagnosis and management. By linking AI methods to concrete clinical needs, this section highlights how advanced imaging analytics can improve diagnostic precision, reduce variability, and ultimately enhance cardiovascular outcomes.
Figure 2: Examples of recent representative AI models for CVD Analysis. Top (Non-vessel). Examples of Networks and Tasks: Convolutional neural network (CNN) [lecun2015deepronneberger2015u], Transformer [vaswani2017attention], Fully convolutional networks (FCN) [long2015fully], U-Net [ronneberger2015u], Autoencoders (AE) and variational autoencoders (VAE) [ehrhardt2022autoencoders], Generative adversarial networks (GAN) [goodfellow2020generative], and Contrastive Language Image Pre-training (CLIP) [zhao2023clip]. Examples of Applications: Calcium Scoring [isgum2012automatic], Image Registration [sheikhjafari2022unsupervised], Whole Heart Segmentation [bui2020simultaneous], Image Synthesis [jaen2022synthetic], Tagged MRI to Cine MRI Transform [liu2021dual], and Report Generation [litjens2019state]. Bottom (Vessel). Examples of Networks and Tasks: Diffusion Model [ho2020denoising], CNN++Recurrent Neural Networks (RNN) [donahue2015long], Tree-structured RNN [lu2023attention], Graph Neural Network (GNN) [wu2020comprehensive], and Reinforcement Learninging (RL) [wang2023deep]. Examples of Applications: Self-supervised Vessel Segmentation [kim2023diffusion], Vessel Stenosis Detection [zreik2018recurrent], Vessel Segmentation [lu2023attention], Vessel Labeling [yang2020cpr] and Guidewire Navigation [kweon2021deep].
3.1 AI Methods
In this section, we review the most widely used AI models for cardiovascular disease (CVD) analysis. An overview of the models and their applications is shown in Figure 2. Convolutional Neural Networks (CNNs) are foundational for image analysis tasks due to their ability to recognize spatial patterns [lecun2015deep]. In CVD analysis, CNNs are used for tasks like heart disease classification, plaque detection in blood vessels, and image classification for disease diagnosis. They excel in analyzing echocardiograms, MRI, and CT scans to identify abnormalities and quantify disease severity. Transformers can analyze time-series data from ECG signals to detect arrhythmias and other heart rhythm disorders [vaswani2017attention]. They are also useful in analyzing 3D imaging data, enabling more accurate analysis of complex cardiac structures by capturing spatial relationships in the heart anatomy. Fully Convolutional Networks (FCNs): FCNs are designed for pixel-level predictions, making them ideal for tasks requiring fine-grained segmentation [long2015fully], such as identifying heart chambers, vessel walls, and plaque deposits in imaging. They are widely used in CVD to delineate cardiac structures and assess the impact of cardiovascular diseases on heart anatomy in MRI and CT scans. UNet is frequently used to segment heart structures [ronneberger2015u], such as the myocardium and ventricles, and to distinguish tissue types, allowing for more accurate assessment of heart health and disease progression in imaging modalities like MRI. Autoencoders (AE) are used for feature extraction, data denoising, and dimensionality reduction [bank2023autoencoders]. In CVD analysis, they are helpful for compressing complex imaging data, enabling more efficient storage and analysis. AEs are also used in unsupervised learning to extract critical features from cardiac images or ECG data, supporting downstream tasks like anomaly detection in imaging or rhythm analysis. Variational Autoencoders (VAEs) extend autoencoders for probabilistic data generation, making them useful for generating synthetic data in CVD research. They can create realistic variations of heart images, which is helpful for augmenting training datasets and improving model robustness. VAEs are particularly useful when working with rare cardiac conditions by generating synthetic examples to balance datasets. Generative Adversarial Networks (GANs) [goodfellow2020generative] are powerful for creating synthetic but realistic images, supporting CVD analysis by augmenting datasets with realistic heart images. GANs can also be used to translate images between modalities, such as generating MRI-like images from CT scans, enhancing data compatibility and model performance when multi-modality data is limited. GANs also play a significant role in medical image segmentation [creswell2018generative, chen2022generative, xun2022generative, xu2020contrast]. Contrastive Language-Image Pretraining (CLIP) aligns images with text descriptions, enabling applications in CVD where imaging findings can be associated with clinical notes [zhao2023clip]. In practice, CLIP allows for cross-modal retrieval, enabling clinicians to search image databases using textual descriptions or retrieve similar cases, streamlining diagnostic workflows and supporting clinical decision-making. More recently, denoising diffusion models, a notable subset of generative models, have recently attracted significant attention in the field of deep learning [yang2023diffusion]. They have demonstrated remarkable utility across a wide range of applications, especially in the enhancement of medical image segmentation [kazerouni2023diffusion]. These models have demonstrated their effectiveness in generating high-quality segmented images, further expanding the possibilities in medical imaging analysis. In CVD, they can generate diverse and realistic images to simulate disease conditions or augment datasets. This is especially valuable for rare diseases or creating data variations that represent disease progression, ultimately improving model training. Recurrent Neural Networks (RNNs) [shi2015convolutional] are designed for sequential data and are applied in CVD analysis for time-series data [azad2019bi] like ECG signals. RNNs are ideal for detecting arrhythmias, analyzing heart rhythm abnormalities, and predicting the progression of heart conditions over time by analyzing patient histories and longitudinal data. Graph Neural Networks (GNNs) [zhou2020graph] are suited for data that can be represented as graphs, making them ideal for modeling the connectivity of heart and vascular structures. In CVD, GNNs can analyze relationships between cardiac segments or vessel pathways [wu2019automated], which supports tasks like vessel segmentation [kong2020learning], detecting aneurysms, and localizing disease within complex vascular networks [gao2020learning]. Reinforcement Learning (RL) [sutton2018reinforcement] is used in CVD for optimizing sequential decision-making, such as developing personalized treatment plans based on predicted patient responses, guidewire navigation [kweon2021deep], and image registration [wang2023deep]. RL can simulate potential treatment outcomes for different interventions, allowing clinicians to tailor therapy plans and improve patient outcomes in areas like heart failure management or post-surgical recovery planning.
Figure 3: Overview of the cardiac image analysis pipeline and examples of structural and functional imaging. Top: Non-vessel. (a): Cardiac segmentation [vesal2020fully], End-diastole (ED) and End-systole (ES) phases detection and Surface Model [meng2022mulvimotion]. (b): Cardiac Motion Tracking [meng2022mulvimotion]. Myocardial blood flow (MBF) maps [jacobs2021automated]. Middle: Medical images serve as inputs for AI models in the CVD diagnosis pipeline. Bottom: Vessel. (c): Segmentation and quantification of vessel structures, with a focus on the Aorta [trentin2015automatic]. Coronary Artery Segmentation [yan2022impact], centerline extraction using segmentation masks [yan2022impact], Coronary Artery Branch Labeling based on centerline structures [yang2020cpr], and detection of Coronary Artery Stenosis [kiricsli2013standardized]. (d): Evaluation of vessel function. Arota flow speed with 3D streamlined visualization [righini2023four]. Interpretation of CT-FFR results showing significant hemodynamic impact [baessato2020incremental], and Coronary Microvascular [clarke2020invasive]. Physiological measurements: CFR = coronary flow reserve; FFR = fractional flow reserve; IMR = index of microvascular resistance.
3.2 AI Pipelines for Cardiac Image Analysis
@@
As illustrated in the left panel of Figure 3, the pipelines start with cardiac image segmentation, which plays a crucial role in numerous clinical applications [zhuang2019evaluation, cetin2023attri]. This process is key to identifying and analyzing essential anatomical structures [peng2016review], as it divides the image into regions with distinct semantic meanings and facilitates the extraction of various quantitative metrics. In the context of non-vessel structures, segmentation aids in calculating important metrics such as the volumes and ejection fractions of the left and right ventricles (LV and RV), atrial sizes, myocardial mass, and wall thickness [chen2020deep]. For vessel structures, segmentation is pivotal for analyzing aspects of coronary arteries [tsai2024uu], including centerline extraction [fu2023prior], plaque identification, and other vascular features. Such detailed segmentation provides invaluable insights for both the diagnosis and treatment planning processes.
@@
As illustrated in the right panel of Figure 3, following segmentation, the AI system engages in feature extraction for precise quantification. This stage involves measuring physical heart parameters like chamber volumes, wall thickness, and myocardial mass [peng2016review]. Utilizing the segmented structures and the features derived, AI models then conduct functional analysis. This encompasses the calculation of ejection fraction, wall motion, and blood flow dynamics, crucial indicators of the heart’s efficiency and overall functioning [peng2016review]. By integrating segmented images, extracted features, and quantification metrics, the AI system forms a comprehensive assessment of cardiac structure and function, invaluable for diagnosing cardiac conditions and formulating treatment plans. The details of methods in Figure 3 are introduced in Section 3.3 for non-vessel structures and Section 3.4 vessel structures. For clinical utility, the AI pipeline presents results in an easily interpretable format, often through 3D models or quantitative maps, compiled into reports for cardiologists’ evaluation. Additionally, it incorporates clinician feedback to refine models, enhancing accuracy and adaptability to new data and insights. This AI-driven approach in cardiology transforms complex imaging data into practical medical information, potentially revolutionizing the diagnosis and treatment of heart diseases.
@@
In the pipeline, the image registration also plays a critical role in analyzing heart structures and diseases, especially when merging data from different imaging modalities or tracking disease progression over time [khalil2018overview, zakeri2022probabilistic, zakeri2023dragnet]. Accurate registration is essential for mapping structural changes precisely and is indispensable in longitudinal studies, treatment planning, and assessing therapeutic efficacy [makela2002review]. The task of cardiac image registration is particularly challenging due to the significant variability in cardiac shape and motion among individuals [toth20183d]. It often involves aligning heart images captured at different times or under varying conditions, such as different phases of the cardiac cycle or using diverse imaging techniques like CT, MRI, or ultrasound [chen2025survey]. Given the dynamic nature of the heart and the influence of respiratory movements, non-rigid registration is crucial for cardiac images [sheikhjafari2022unsupervised]. This approach is key to understanding disease progression or treatment effects by comparing temporal image sequences. Advancements in AI models are vital for enhancing registration accuracy and efficiency. The development of real-time registration systems is particularly significant for their potential application in surgical or interventional settings [wang2023deep], promising to significantly improve patient outcomes and procedural precision in cardiac care [lara2022deep].